# Newton numbers, vanishing polytopes and algebraic degrees

### v1

@@ 6.4.2 Negligible polytopes are $B_{k}$-polytopes
Consider a subset $\mathbf{\Delta}\subset\mathbf{P}$ consisting of vertices of an $r$-dimensional simplex. For a pair of opposite faces $F_{1}\sqcup F_{2}=\mathbf{\Delta}$, consider the stripe parallel to $\text{aff}(F_{1}-F_{2})$ such that $F_{1}$ and $F_{2}$ are on its boundary. Then for any point $X$ onside such $2^{r}$ stripes the union $\mathbf{P}\cup X$ has width greater than one. ∎
@@
Consider a generic lattice point $X\in\mathbb{Z}^{n-m}$. By Lemma 5.27 we have the inequality $\nu^{e}_{X}(\mathbf{P})\geq\nu_{C}(P)$, so $\nu^{e}_{X}(\mathbf{P})=0$. Note that $\nu^{e}_{X}(\mathbf{P})=\nu^{e}_{X}(\mathbf{P}\cup X)$ (it is implied e.g. by Remark 5.21). Denote by $\mathbf{A}\supset O$ the set $(\mathbf{P}\cup X)-X$, i.e. the set $\mathbf{P}\cup X$ shifted by the vector $-X$. By Lemma 6.2 it is dual defective.
Denote by $\mathbb{Z}_{\mathbf{A}}^{n}$ the minimal sublattice of $\mathbb{Z}^{n}$ containing $\mathbf{A}$. From the Furukawa-Ito classification Theorem 6.4 (and Remark 6.5) we obtain that there are integers $k>c$, two projections $\pi_{k}:\mathbb{Z}_{\mathbf{A}}^{n}\to\mathbb{Z}^{k}$ and $p:\mathbb{Z}_{\mathbf{A}}^{n}/\mathbb{Z}^{k}\to\mathbb{Z}^{n-k-c}$ such that:
@@
    $\pi_{k}(\mathbf{A})=Conv(O,q_{1},\dots,q_{k})$ is the unit simplex in $\mathbb{Z}^{k}$
@@
Without loss of generality we can assume that $\pi_{k}(O)=O$, where $O$ is the origin (otherwise we can compose $\pi$ with an appropriate affine automorphism of $\mathbb{Z}^{k}$). Consider the affine extension of $\pi_{k}$ on $\mathbb{R}^{n}$ and denote it also by $\pi_{k}$. Note that the image of the lattice $\pi_{k}(\mathbb{Z}^{n})$ may not be contained in the lattice $\mathbb{Z}^{k}$.
@@
By Proposition 6.30 we have $\pi_{k}(\mathbb{R}^{n-m})=O$.
@@
There are positive integers $a_{i}^{j}\in\mathbb{Z}_{>0}^{n}$ and a renumbering of the standard basis $e_{i}^{(j)}$ of $\mathbb{Z}^{n}$ so that the map $\pi_{k}$ can be written in the following way:
@@
By Remark 6.32 we have $\pi_{k}(\mathbb{R}^{n-m})=O$, so the first $n-m$ elements $e_{0}^{(1)},\dots,e_{0}^{(n-m)}$ are the standard basis of $\mathbb{Z}^{n-m}$. For each standard basis vector $v_{i}$ of $\mathbb{Z}^{m}_{\geq 0}$ there exists $a_{i}\in\mathbb{Z}_{>0}$ and a point of $\mathbf{A}$ whose projection along $\mathbb{Z}^{n-m}$ is $a_{i}v_{i}$, since $\operatorname{Conv}(\mathbf{A})$ is convenient in $C$. So $\pi_{k}(a_{i}v_{i})$ is either the origin $O$ or a basis vector $q_{j}$. Considering this way all the coordinate rays of $\mathbb{Z}^{m}_{\geq 0}$, we obtain the required form. ∎
@@
First let us prove the $B_{k}$-theorem if all these points have unit coordinate.
@@
Let us prove that instead of $\pi_{k}$ we can consider the following coordinate projection $\widetilde{\pi}_{k}$ (with the same $p$) so that the conditions of the Furukawa-Ito Theorem 6.4 are preserved:
@@
Note that for any point $x\in\mathbf{A}$, at most one of coordinates $e_{\alpha}^{(\beta)},\alpha\geq 1$ can be non-zero, and the non-zero coordinate must be equal to $1$. Denote the sets $\widetilde{\mathbf{A}}_{i}\subset\mathbb{Z}^{n}/\mathbb{Z}^{k}$ the same way as $\mathbf{A}_{i}$ but for projection $\widetilde{\pi}_{k}$. Denote the sets
@@
Denote by $E_{0}$ the lattice spanned by $\{e_{0}^{(i)}/\mathbb{Z}^{k},i=1,\dots,j_{0}\}$. Note that $\langle\mathbf{A}_{0}-\mathbf{A}_{0}\rangle=E_{0}$ since $\operatorname{Conv}(\mathbf{A})$ is convenient, and that $\langle\mathbf{A}_{\alpha}^{(\beta)}-\mathbf{A}_{\alpha}^{(\beta)}\rangle\subset E_{0}$ for any $1\leq\alpha\leq k$ and $1\leq\beta\leq j_{\alpha}$. Recall that $p(\mathbf{A}_{0})\ast\dots\ast p(\mathbf{A}_{k})$ is of join type. So $p(\langle\mathbf{A}_{\alpha}^{(\beta)}-\mathbf{A}_{\alpha}^{(\beta)}\rangle)=0$ for any $1\leq\alpha\leq k$ and $1\leq\beta\leq j_{\alpha}$. Thus $p(\widetilde{\mathbf{A}}_{0})\ast\dots\ast p(\widetilde{\mathbf{A}}_{k})$ is also of join type since $p(\widetilde{\mathbf{A}}_{\alpha})=0$ for all $\alpha=1,\dots,k$. ∎
In what follows, we prove the $B_{k}$-theorem for arbitrary $a_{i}^{(j)}\in\mathbb{Z}_{\geq 0}$ by induction on the dimension $n$.
@@
If all the $a_{i}^{(j)}$ are equal to $1$, then this case is already proved in Lemma 6.35. Otherwise, by the induction hypothesis, the polytope $\operatorname{Conv}(\mathbf{A})\cap C^{(1)}$ is a $B_{k}$-polytope. Note that (here we essentially use that the corresponding $a_{i}^{(j)}$ are equal to $1$) the coordinate projection of $\mathbf{A}\setminus C^{(1)}$ to $C^{(1)}$ is only the origin $O$. Then the polytope $\operatorname{Conv}(\mathbf{A})$ is also a $B_{k}$-polytope (different from $\operatorname{Conv}(\mathbf{A})\cap C^{(1)}$ only in the first summand). ∎
@@
Consider all pairs $i,j$ such that $a_{i}^{(j)}\geq 2$. Consider any points $D_{i}^{(j)}$ such that $\pi_{k}^{O}(a_{i}^{(j)}e_{i}^{(j)})=\pi_{k}^{O}(D_{i}^{(j)})$ if $i\neq j$; such points exist since $\operatorname{Conv}(\mathbf{A})$ is convenient. Consider the set $\mathbf{D}=\bigcup_{i,j}D_{i}^{(j)}\cup(\mathbf{A}\cap C^{(1)})$ Note that $\operatorname{Conv}(\mathbf{D})$ is convenient in the cone $C$ and
@@
Recall that the Newton number is monotonous (see Remark 5.29). Thus, since $\operatorname{Conv}(\mathbf{A})$ is negligible, the convex hull $\operatorname{Conv}(\mathbf{D})$ is also negligible in $C$, hence $\operatorname{Conv}(\mathbf{A})\cap C^{(1)}$ is negligible in $C^{(1)}$.

### v2

@@ 6.4.2 Negligible polytopes are $\mathbf{B_{k}}$-polytopes
Consider a subset $\mathbf{\Delta}\subset\mathbf{P}$ consisting of the vertices of an $r$-dimensional simplex. For a pair of opposite faces $F_{1}\sqcup F_{2}=\mathbf{\Delta}$, consider the stripe parallel to $\text{aff}(F_{1}-F_{2})$ such that $F_{1}$ and $F_{2}$ are on its boundary. Then for any point $X$ outside these $2^{r}$ stripes, the union $\mathbf{P}\cup X$ has width greater than $1$. ∎
@@
Consider a generic lattice point $X\in\mathbb{Z}^{n-m}$. By Lemma 5.28, we have the inequality $\nu^{e}_{X}(\mathbf{P})\geq\nu_{C}(P)$, so $\nu^{e}_{X}(\mathbf{P})=0$. Note that $\nu^{e}_{X}(\mathbf{P})=\nu^{e}_{X}(\mathbf{P}\cup X)$ (it is implied e.g. by Remark 5.22). Denote by $\mathbf{A}\supset O$ the set $(\mathbf{P}\cup X)-X$, i.e. the set $\mathbf{P}\cup X$ shifted by the vector X-X. By Lemma 6.3 it is dual defective.
Denote by $\mathbb{Z}_{\mathbf{A}}^{n}$ the minimal affine sublattice of $\mathbb{Z}^{n}$ containing $\mathbf{A}$. From the Furukawa-Ito classification Theorem 6.5 (and Remark 6.6), we obtain that there are integers $k>c$, and two projections $\pi_{k}:\mathbb{Z}_{\mathbf{A}}^{n}\to\mathbb{Z}^{k}$ and $p:\mathbb{Z}_{\mathbf{A}}^{n}/\mathbb{Z}^{k}\to\mathbb{Z}^{n-k-c}$ such that:
@@
    $\pi_{k}(\mathbf{A})=Conv(O,q_{1},\dots,q_{k})$ is the unit simplex in $\mathbb{Z}^{k}$.
@@
Without loss of generality, we can assume that $\pi_{k}(O)=O$, where $O$ is the origin (otherwise we can compose $\pi$ with an appropriate affine automorphism of $\mathbb{Z}^{k}$). Consider the affine extension of $\pi_{k}$ on $\mathbb{R}^{n}$ and denote it also by $\pi_{k}$. Note that the image of the lattice $\pi_{k}(\mathbb{Z}^{n})$ may not be contained in the lattice $\mathbb{Z}^{k}$.
@@
By Proposition 6.36, we have $\pi_{k}(\mathbb{R}^{n-m})=O$.
@@
There are positive integers $a_{i}^{(j)}\in\mathbb{Z}_{>0}$ and a renumbering of the standard basis $e_{i}^{(j)}$ of $\mathbb{Z}^{n}$ such that the map $\pi_{k}$ can be written as follows:
@@
By Remark 6.38, we have $\pi_{k}(\mathbb{R}^{n-m})=O$, so the first nmn-m elements $e_{0}^{(1)},\dots,e_{0}^{(n-m)}$ are the standard basis of $\mathbb{Z}^{n-m}$. For each standard basis vector $v_{i}$ of $\mathbb{Z}^{m}_{\geq 0}$, there exists $a_{i}\in\mathbb{Z}_{>0}$ and a point in $\mathbf{A}$ whose projection along $\mathbb{Z}^{n-m}$ is $a_{i}v_{i}$, since $\operatorname{Conv}(\mathbf{A})$ is convenient in $C$. So $\pi_{k}(a_{i}v_{i})$ is either the origin $O$ or a basis vector $q_{j}$. Considering this way all the coordinate rays of $\mathbb{Z}^{m}_{\geq 0}$, we obtain the required form. ∎
@@
First, let us prove the $B_{k}$-theorem if all these points have unit coordinates.
@@
Let us prove that instead of $\pi_{k}$, we can consider the coordinate projection $\widetilde{\pi}_{k}$ (with the same $p$) such that the conditions of the Furukawa-Ito Theorem 6.5 are preserved:
@@
Note that for any point $x\in\mathbf{A}$, at most one of the coordinates $e_{\alpha}^{(\beta)}$ with $\alpha\geq 1$ can be nonzero, and any such nonzero coordinate must be equal to $1$. Define the sets $\widetilde{\mathbf{A}}_{i}\subset\mathbb{Z}^{n}/\mathbb{Z}^{k}$ analogously to $\mathbf{A}_{i}$, but using the projection $\widetilde{\pi}_{k}$. Denote the sets
@@
Denote by $E_{0}$ the lattice spanned by $\{e_{0}^{(i)}/\mathbb{Z}^{k},i=1,\dots,j_{0}\}$. Note that $\langle\mathbf{A}_{0}-\mathbf{A}_{0}\rangle=E_{0}$ since $\operatorname{Conv}(\mathbf{A})$ is convenient, and that $\langle\mathbf{A}_{\alpha}^{(\beta)}-\mathbf{A}_{\alpha}^{(\beta)}\rangle\subset E_{0}$ for any $1\leq\alpha\leq k$ and $1\leq\beta\leq j_{\alpha}$. Recall that $p(\mathbf{A}_{0})\ast\dots\ast p(\mathbf{A}_{k})$ is of join type. Therefore, $p(\langle\mathbf{A}_{\alpha}^{(\beta)}-\mathbf{A}_{\alpha}^{(\beta)}\rangle)=0$ for any $1\leq\alpha\leq k$ and $1\leq\beta\leq j_{\alpha}$. Thus $p(\widetilde{\mathbf{A}}_{0})\ast\dots\ast p(\widetilde{\mathbf{A}}_{k})$ is also of join type since $p(\widetilde{\mathbf{A}}_{\alpha})=0$ for all $\alpha=1,\dots,k$. ∎
In what follows, we prove the $B_{k}$-theorem for arbitrary $a_{i}^{(j)}\in\mathbb{Z}_{>0}$ by induction on the dimension $n$.
@@
If all the $a_{i}^{(j)}$ are equal to $1$, then this case is already proved in Lemma 6.41. Otherwise, by the induction hypothesis, the polytope $\operatorname{Conv}(\mathbf{A})\cap C^{(1)}$ is a $B_{k}$-polytope. Note that (here we essentially use that the corresponding $a_{i}^{(j)}$ are equal to $1$) the coordinate projection of $\mathbf{A}\setminus C^{(1)}$ to $C^{(1)}$ is only the origin $O$. Then the polytope $\operatorname{Conv}(\mathbf{A})$ is also a $B_{k}$-polytope (differing from $\operatorname{Conv}(\mathbf{A})\cap C^{(1)}$ only in the first summand). ∎
@@
Consider all pairs $i,j$ such that $a_{i}^{(j)}\geq 2$. Consider points $D_{i}^{(j)}$ such that $\pi_{k}^{O}(a_{i}^{(j)}e_{i}^{(j)})=\pi_{k}^{O}(D_{i}^{(j)})$ when $i\neq j$; such points exist since $\operatorname{Conv}(\mathbf{A})$ is convenient. Consider the set $\mathbf{D}=\bigcup_{i,j}D_{i}^{(j)}\cup(\mathbf{A}\cap C^{(1)})$ . Note that $\operatorname{Conv}(\mathbf{D})$ is convenient in the cone $C$ and
@@
Recall that the Newton number is monotonic (see Remark 5.30). Thus, since $\operatorname{Conv}(\mathbf{A})$ is negligible, the convex hull $\operatorname{Conv}(\mathbf{D})$ is also negligible in $C$. Therefore, $\operatorname{Conv}(\mathbf{A})\cap C^{(1)}$ is negligible in $C^{(1)}$.